The height (in centimeters) of a candle is a linear function of the amount of time (in hours) it has been burning, when graphed, the function gives a line with a slope of -0.6. Suppose that the height of the candle after 20 hrs is 12 centimeters. What was the height of the candle after 15 hrs?
Answer by ankor@dixie-net.com(22740)

The height (in centimeters) of a candle is a linear function of the amount of
time (in hours) it has been burning, when graphed, the function gives a line
with a slope of -0.6.
Suppose that the height of the candle after 20 hrs is 12 centimeters.
:
Use the point/slope equation: y - y1 = m(x - x1)
where m =-.6; x = 20; y = 12
:
y - 12 = -.6(x - 20)
y - 12 = -.6x + 12
y = -.6x + 12 + 12
y = -.6x + 24, is the equation (y = length after x hours)
:
"What was the height of the candle after 15 hrs?"
y = -.6(15) + 24
y = -9 + 24
y = 15 inches after 15 hrs